SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jun 7, 2005)

Operations

  • On the 4th and 5th of June, a problem occurred in the old TSF system but didn't affect data quality. See more details under DCT/DCZ discussion below and a posting at the Trigger Operation HN.
  • The Trigger monthly status report will be given tomorrow after the detector operation meeting ~4pm.

EMT

  • Dan showed some new EMT fastmon plots demonstrating the turn-on's using mu-pair events. Note that the binormal errors in the plots aren't handled properly in PAW, and may have to calculate by hand based on Kris' experience. After finalizing these plots, they will be available in live fastmon .ps archive.
  • Dan mentioned his fastmon job crashed when it processed more than 35K events. Rainer suggested that there might be a potential memory leak and adviced Dan to post it to the online event processing HN.
  • L3 requires EMC clusters in selecting mu-pair, so the L3 mu-pair skims won't have a pure DCH mu-pair sample for a study of M efficiency. The bhabha flag in OEP input line was also briefly mentioned since Dan wants to run L1 fastmon to see bhabha energy peak for all bhabha events.

DCT/DCZ

  • The Old TSF occupancy plots had some dips in the TSF-X14 and TSF-X15 in two separate occasions. In both cases, the DQM shifter noticed that the colored histograms in the NEW vs OLD TSF comparison plots which indicated there were differences in pattern, timing, and location of the TSF segments in the two systems. (see plots for run 54988 and 55000). Su Dong mentioned that this might not be a new problem since similar things might have happened last June/July. Back then, there wasn't the new vs old TSF comparison plots to help investigate, so he wasn't sure if the problem was with the DCH or the DCT. The location of the problem seems very random, and the problem went away after reconfiguring the system. The short term instruction for Navigators is updated to include instruction regarding this problem. We want to keep a record of this problem and hope to have a better understanding of the problem.
  • Rainer reminded us that we still need to make a decision about damage flag for the ZPD and BLT. Currently, there are no damage flag for ZPD and BLT. A few weeks ago, there were some runs with ZPD damages that affected about 0.3% of the events. It was small enough fraction that shouldn't bias trigger decision, so the data are principle good for physics still. Perhaps we could come up with some sort of decision about this in our next trigger operation meeting.
  • Jamie reminded us that we will need to rebuild the ZPD ROM library for the next online release built that based on 18.1.0 or higher at IR2. More info can be seen in the trigger software HN.

L3

  • Online release 13.1.0 was installed at IR2 today, and Rainer got a page about L3 showing 100% damage when the new release was in. It turned out that COSMICs_NODCH configuration was used but they need to take DCT and DCZ out of the partitions.

Data Quality and Monitoring

  • Kris were looking at Run 1 reprocessing. Questions were raised about L3 EMC occupancy having only ~15% coverage in some early runs in Run 1. Rainer remembered that the L3 EMC cluster were in after Jan 2000 or so. See Rainer's web page for filter configuration revision for Run1-5.
  • Kris needed to make a new tag of the fastMon for Jim Hamilton to build a new release. Selina is working on the last stage of testing the DCZ damage check using JAS interface. The L1TFMon code is ready to be checked in.
  • OepSequences in JAS should have the BLT/PTD replaced with the BLT/ZPD plots. It turned out that we haven't been monitoring the DCZ in the "DataFlow/OEP" plots of JAS all this time. These plots would provide basic event structure with information such as number of bytes per ROM, module IDs of ROMs reporting data, etc.. Last Friday, we spent about 15 minutes of the L3 meeting for discussions related to this. It should be very easy to implement.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct the above items, specially for those who attended the meeting. Thanks.